# Thunder: Harnessing Unused GPU Cycles in Large-Scale Training via Opportunistic Execution

### v2

@@ Abstract
In large-scale deep learning (DL) tasks, such as those involving modern language models, training inefficiencies frequently lead to periods of idle GPU capacity. This paper introduces Thunder, an approach that co-locates additional inference tasks alongside primary training jobs. By monitoring runtime behavior, it adaptively detects underutilized GPU intervals and fills them with online or offline inference workloads. Evaluations under mainstream distributed training regimes show that Thunder can significantly boost GPU utilization, achieving up to a 14× increase in offline inference throughput compared to traditional scheduling methods, while also reducing the 95th-percentile latency of online inference by 67% relative to MPS—without sacrificing the throughput of the main training process.
@@
Distributed Training GPU Resource Optimization Opportunistic Execution
@@ 1 Introduction
In this paper, we present Thunder, a system that leverages an opportunistic inference execution mechanism, to exploit idling GPU resources of distributed training and increase aggregated throughputs of GPUs. First, it allows inference instances to collocate with training instances according to their memory demands and GPU idling characteristics. Second, it adopts a Bubble Monitor to detect GPU idling timing in real time. Third, it builds a CUDA Kernel Scheduler to issue tokens to collocated inference instances, in which the Kernel Barrier decides to release inference CUDA kernels to fill training bubbles. In summary, our contributions are as follows:
@@
  • We build and evaluate the Thunder system. The experiments show that Thunder significantly improves GPU utilization of various distributed training modes, delivering additional up to 14$\times$ offline inference throughputs than TGS and 67% reduction in online inference p95 latency than MPS, while guaranteeing training throughput.
@@ 3 System Design
Figure 2: The system architecture of Thunder.
@@ 3.1 Overview
Thunder builds on the key idea of speculatively filling inference workloads into bubble periods [^1] For simplicity, we refer to the GPU-idling period in training mentioned above as bubbles. produced by distributed training tasks, as illustrated in Figure 2. It consists of a kernel control layer and a coordination layer.
At the kernel control layer, training and inference workloads can be observed and manipulated in the view of CUDA kernels. First, Thunder collocates one or more additional inference instances with a training instance to share a GPU, preparing the runtime context to speculatively fill inference workloads. Second, to observe the GPU utilization of training instances, Thunder initiates the Bubble Monitor (BM) for each training worker to detect idle periods. It is suitable for general communication-optimized distributed DL frameworks, such as DeepSpeed [13], Megatron-LM [21]. BM intercepts❶ and counts❸ the CUDA kernels, issued at runtime by training instances, which are dispatched to GPU for execution❷. Third, Thunder leverages the Kernel Barrier (KB) to take charge of either blocking or forwarding CUDA kernels of inference instances to the GPU.
At the coordination layer, the CUDA Kernel Scheduler (CKS) on each node is responsible for recognizing training bubbles❹ and deciding speculative inference execution timings of collocated inference instances❺. Thunder introduces two distinct mechanisms tailored for offline and online inference tasks, acknowledging their varying degrees of SLO-sensitivity (Service Level Objective), which is explained in Section 3.3.3.
@@ 3.2 Collocation Basics
Before running the speculative filling workflow, collocation should be done in advance. There are several factors explicitly affecting the collocation policies, including distinct spatial and temporal GPU requirements of tasks (as shown in Section 2.1), risks of OOM errors, training interruptions and throughput degradation due to severe resource contention. According to profilings, Thunder adheres to the following principles:
  • Principle-I: The sum of GPU memory occupied at the peak of each collocated instance must be less than the upper limit (e.g., 40GB) of a specific GPU. Based on it, Thunder collocates inference instances as much as possible to enhance aggregated throughput and GPU utilization.
@@ 3.3 Speculative Inference Filling
Thunder employs a hijacking statistics method. Specifically, it mounts the monitoring library (MLib) to training instances. The MLib intercepts the CUDA kernels and records the counts periodically (e.g., 2ms), which are saved to a sliding window maintained by the BM. Concurrently, the BM calculates the number of continuous zero-counts of sliding windows and then sends it to the CUDA Kernel Scheduler on the node.
@@
For online inference workloads, KB additionally adopts a real-time pull-and-execute mechanism to meet the demands of stringent SLOs, since it is challenging to fill online requests into training bubbles without mutual interference due to the unpredictability of workloads. In this case, KB proactively pulls online requests one by one from the request queue, upon receiving the idle signal (set by Algorithm 1). Based on Principle-II in Section 3.2, Thunder ensures to handle at least one request during bubbles of each training iteration. Moreover, to avoid training resuming immediately after pulling one request, CKS preemptively sets the status to busy, according to profiling information on training iteration time (e.g., 1.5 seconds). During the busy status, requests are handled by other inference instances.
@@ 4 Implementation
We implement a prototype of Thunder with 2k C LOCs and 2k Python LOCs for evaluations. In the prototype, all training and inference tasks run with NVIDIA-docker runtime.
@@ 5.1 Methodology
Experiment testbed. We evaluate Thunder on a GPU server with 4 * NVIDIA A100-40GB, equipped with PyTorch v1.11, DeepSpeed v0.11.1, CUDA v11.7.
@@
Baselines. We compare Thunder with the following methods.
@@ 5.2 Speculative Inference Filling Performance
Offline Inference Filling. Figures LABEL:fig:ddp-offline,LABEL:fig:mp-offline and LABEL:fig:pp-offline demonstrate that Thunder delivers high throughputs for offline inferences with training throughput guarantees in DP, MP and PP modes. Considering the primary training workload, all baselines, except Co-Exec, generally maintain the performance. For the collocated offline inference workloads, Thunder provides 23-84% throughput of Exclusive and 33-94% of Co-Exec, best in other baselines. However, the Exclusive requires one additional GPU, and Co-Exec significantly reduces collocated training throughput (e.g., up to 28% in the RoBERTa-Resnet case), failing to meet the goals in Section 2.2.
In DP cases, inference throughput is 1.23$\times$-3.5$\times$ of MPS, and 2.9$\times$-14$\times$ of TGS. The underperformance of TGS is mainly due to its inadequate bubble detection and relatively conservative time-share strategy, while MPS statically limits GPU resources available to the inference process. In MP cases, Thunder achieves the highest aggregated throughputs in the first three cases with training throughput guarantees. Specifically, the inference throughput increases by up to 80% and 11$\times$ compared to MPS and TGS, respectively. In PP cases, the advantages of Thunder become relatively marginal to DP and MP cases. Co-Exec achieves significantly higher inference throughput but sacrifices 3%-9% of training throughput. Thunder’s performance is comparable to MPS and superior to TGS. The underlying reason is that though dividing a mini-batch into small micro-batches shortens bubbles, it leaves GPU consistently underutilized, thus allowing it to sufficiently handle inference workloads in Co-Exec mode. In the future, for the PP scenario, we aim to execute inference workloads concurrently as much as possible until the BM observes that training iterations have lengthened.
Online Inference Filling. Figures  LABEL:fig:ddp-online,LABEL:fig:mp-online,LABEL:fig:pp-online demonstrate that Thunder can deliver low p95 latencies of online inference with training throughput guarantees. Thunder consistently shows the lowest p95 latency of inferences, trailing only behind the Exclusive mode, while maintaining standard training throughputs. This advantage is owing to the proactive pull-and-execute mechanism. In DP cases, shown in Figure LABEL:fig:ddp-online, Thunder reduces p95 by up to 61% and 67% compared to Co-Exec and MPS, respectively. In MP cases, it lowers p95 by an average of 40% compared to Co-Exec and 33% compared to MPS. Similar to offline cases above, the gains in PP modes diminish, but Thunder still maintains the best tail latency performance except Exclusive.
@@ 5.3 Multi-instance Support
Aggregated Throughput Improvement with Multiple Inference Instances Support. As Principle-I in Section 3.2 mentions, Thunder supports collocating multiple inference instances to enhance GPU utilization. Figure LABEL:fig:varysize-performance shows that Thunder achieves a sub-linear growth in inference throughput while ensuring training throughput, with increasing collocated inference instances. In DP cases depicted in Figure LABEL:fig:ddp-varysize, Thunder outperforms Exclusive by achieving an additional 35%-123% in inference throughput when the number of instances ranges from 1 to 4, with a maximum training throughput reduction of less than 7%. Although Co-Exec surpasses all other baselines in aggregated throughput, it leads to a substantial degradation in training, up to 61%. Consistent trends are observed in the MP scenario shown in Figure LABEL:fig:mp-varysize. Specifically, Thunder matches the offline throughput performance of Co-Exec while avoiding the latter’s detrimental impact on training, which can reach up to 22%. Notably, with 4 instances, the inference throughput soars to 299% more than that of Exclusive.
@@ 5.4 System Overhead
Figure 4: The system overhead of Thunder.
@@ 6 Related Work
GPU Sharing. The straightforward method to enhance GPU utilization effectively is to share a single GPU with multiple DL tasks. Existing works can be categorized into temporal and spatial sharing approaches. For temporal sharing, Antman [24] implicitly inserts training jobs during iterations. Based on this, TGS [23] provides high transparency. However, both methods are not suitable for LLMs due to large memory footprints. As for spatial sharing, MIG [16] supports the physical isolation of GPU devices but lacks adaptability. MPS [17] is widely used in DL systems [5, 25], while it can not dynamically consume idle compute resources completely due to static allocation. Aiming at distributed training, Thunder detects bubbles timely and speculatively fills inference workloads to improve GPU utilization.
@@ 7 Conclusion
Nowadays, Deep Learning revolutionizes various aspects of life. However, GPUs used for training these DL applications are usually underutilized, yielding massive compute and memory fragments. We observe that moderate inference workloads are well-suited to fill up these GPU fragmentations. In this paper, we present Thunder, which collocates distributed training with online/offline inference instances, to speculatively serve inference workloads, significantly improving GPU utilization. The results show that Thunder can exploit the idling GPU resources in various distributed training modes, delivering additional up to 14$\times$ offline inference throughputs than TGS and 67% reduction in online inference p95 latency than MPS, while guaranteeing collocated training throughput.

### v3

# SpecInF: Exploiting Idle GPU Resources in Distributed DL Training via Speculative Inference Filling

@@ Abstract
Deep Learning (DL), especially with Large Language Models (LLMs), brings benefits to various areas. However, DL training systems usually yield prominent idling GPU resources due to many factors, such as resource allocation and collective communication. To improve GPU utilization, we present SpecInF, which adopts a Speculative Inference Filling method to exploit idle GPU resources. It collocates each primary training instance with additional inference instances on the same GPU, detects the training bubbles and adaptively fills with online or offline inference workloads. Our results show that SpecInF can effectively enhance GPU utilization under mainstream parallel training modes, delivering additional up to 14$\times$ offline inference throughputs than TGS and 67% reduction in online inference p95 latency than MPS, while guaranteeing collocated training throughput.
@@
Distributed Training Collocation Speculative Inference Filling
@@ 1 Introduction
In this paper, we present SpecInF, a system that leverages a Speculative Inference Filling mechanism, to exploit idling GPU resources of distributed training and increase aggregated throughputs of GPUs. First, it allows inference instances to collocate with training instances according to their memory demands and GPU idling characteristics. Second, it adopts a Bubble Monitor to detect GPU idling timing in real time. Third, it builds a CUDA Kernel Scheduler to issue tokens to collocated inference instances, in which the Kernel Barrier decides to release inference CUDA kernels to fill training bubbles. In summary, our contributions are as follows:
@@
  • We build and evaluate the SpecInF system. The experiments show that SpecInF significantly improves GPU utilization of various distributed training modes, delivering additional up to 14$\times$ offline inference throughputs than TGS and 67% reduction in online inference p95 latency than MPS, while guaranteeing training throughput.
@@ 3 System Design
Figure 2: The system architecture of SpecInF.
@@ 3.1 Overview
SpecInF builds on the key idea of speculatively filling inference workloads into bubble periods [^1] For simplicity, we refer to the GPU-idling period in training mentioned above as bubbles. produced by distributed training tasks, as illustrated in Figure 2. It consists of a kernel control layer and a coordination layer.
At the kernel control layer, training and inference workloads can be observed and manipulated in the view of CUDA kernels. First, SpecInF collocates one or more additional inference instances with a training instance to share a GPU, preparing the runtime context to speculatively fill inference workloads. Second, to observe the GPU utilization of training instances, SpecInF initiates the Bubble Monitor (BM) for each training worker to detect idle periods. It is suitable for general communication-optimized distributed DL frameworks, such as DeepSpeed [13], Megatron-LM [21]. BM intercepts❶ and counts❸ the CUDA kernels, issued at runtime by training instances, which are dispatched to GPU for execution❷. Third, SpecInF leverages the Kernel Barrier (KB) to take charge of either blocking or forwarding CUDA kernels of inference instances to the GPU.
At the coordination layer, the CUDA Kernel Scheduler (CKS) on each node is responsible for recognizing training bubbles❹ and deciding speculative inference execution timings of collocated inference instances❺. SpecInF introduces two distinct mechanisms tailored for offline and online inference tasks, acknowledging their varying degrees of SLO-sensitivity (Service Level Objective), which is explained in Section 3.3.3.
@@ 3.2 Collocation Basics
Before running the speculative filling workflow, collocation should be done in advance. There are several factors explicitly affecting the collocation policies, including distinct spatial and temporal GPU requirements of tasks (as shown in Section 2.1), risks of OOM errors, training interruptions and throughput degradation due to severe resource contention. According to profilings, SpecInF adheres to the following principles:
  • Principle-I: The sum of GPU memory occupied at the peak of each collocated instance must be less than the upper limit (e.g., 40GB) of a specific GPU. Based on it, SpecInF collocates inference instances as much as possible to enhance aggregated throughput and GPU utilization.
@@ 3.3 Speculative Inference Filling
SpecInF employs a hijacking statistics method. Specifically, it mounts the monitoring library (MLib) to training instances. The MLib intercepts the CUDA kernels and records the counts periodically (e.g., 2ms), which are saved to a sliding window maintained by the BM. Concurrently, the BM calculates the number of continuous zero-counts of sliding windows and then sends it to the CUDA Kernel Scheduler on the node.
@@
For online inference workloads, KB additionally adopts a real-time pull-and-execute mechanism to meet the demands of stringent SLOs, since it is challenging to fill online requests into training bubbles without mutual interference due to the unpredictability of workloads. In this case, KB proactively pulls online requests one by one from the request queue, upon receiving the idle signal (set by Algorithm 1). Based on Principle-II in Section 3.2, SpecInF ensures to handle at least one request during bubbles of each training iteration. Moreover, to avoid training resuming immediately after pulling one request, CKS preemptively sets the status to busy, according to profiling information on training iteration time (e.g., 1.5 seconds). During the busy status, requests are handled by other inference instances.
@@ 4 Implementation
We implement a prototype of SpecInF with 2k C LOCs and 2k Python LOCs for evaluations. In the prototype, all training and inference tasks run with NVIDIA-docker runtime.
@@ 5.1 Methodology
Experiment testbed. We evaluate SpecInF on a GPU server with 4 * NVIDIA A100-40GB, equipped with PyTorch v1.11, DeepSpeed v0.11.1, CUDA v11.7.
@@
Baselines. We compare SpecInF with the following methods.
@@ 5.2 Speculative Inference Filling Performance
Offline Inference Filling. Figures LABEL:fig:ddp-offline,LABEL:fig:mp-offline and LABEL:fig:pp-offline demonstrate that SpecInF delivers high throughputs for offline inferences with training throughput guarantees in DP, MP and PP modes. Considering the primary training workload, all baselines, except Co-Exec, generally maintain the performance. For the collocated offline inference workloads, SpecInF provides 23-84% throughput of Exclusive and 33-94% of Co-Exec, best in other baselines. However, the Exclusive requires one additional GPU, and Co-Exec significantly reduces collocated training throughput (e.g., up to 28% in the RoBERTa-Resnet case), failing to meet the goals in Section 2.2.
In DP cases, inference throughput is 1.23$\times$-3.5$\times$ of MPS, and 2.9$\times$-14$\times$ of TGS. The underperformance of TGS is mainly due to its inadequate bubble detection and relatively conservative time-share strategy, while MPS statically limits GPU resources available to the inference process. In MP cases, SpecInF achieves the highest aggregated throughputs in the first three cases with training throughput guarantees. Specifically, the inference throughput increases by up to 80% and 11$\times$ compared to MPS and TGS, respectively. In PP cases, the advantages of SpecInF become relatively marginal to DP and MP cases. Co-Exec achieves significantly higher inference throughput but sacrifices 3%-9% of training throughput. SpecInF’s performance is comparable to MPS and superior to TGS. The underlying reason is that though dividing a mini-batch into small micro-batches shortens bubbles, it leaves GPU consistently underutilized, thus allowing it to sufficiently handle inference workloads in Co-Exec mode. In the future, for the PP scenario, we aim to execute inference workloads concurrently as much as possible until the BM observes that training iterations have lengthened.
Online Inference Filling. Figures  LABEL:fig:ddp-online,LABEL:fig:mp-online,LABEL:fig:pp-online demonstrate that SpecInF can deliver low p95 latencies of online inference with training throughput guarantees. SpecInF consistently shows the lowest p95 latency of inferences, trailing only behind the Exclusive mode, while maintaining standard training throughputs. This advantage is owing to the proactive pull-and-execute mechanism. In DP cases, shown in Figure LABEL:fig:ddp-online, SpecInF reduces p95 by up to 61% and 67% compared to Co-Exec and MPS, respectively. In MP cases, it lowers p95 by an average of 40% compared to Co-Exec and 33% compared to MPS. Similar to offline cases above, the gains in PP modes diminish, but SpecInF still maintains the best tail latency performance except Exclusive.
@@ 5.3 Multi-instance Support
Aggregated Throughput Improvement with Multiple Inference Instances Support. As Principle-I in Section 3.2 mentions, SpecInF supports collocating multiple inference instances to enhance GPU utilization. Figure LABEL:fig:varysize-performance shows that SpecInF achieves a sub-linear growth in inference throughput while ensuring training throughput, with increasing collocated inference instances. In DP cases depicted in Figure LABEL:fig:ddp-varysize, SpecInF outperforms Exclusive by achieving an additional 35%-123% in inference throughput when the number of instances ranges from 1 to 4, with a maximum training throughput reduction of less than 7%. Although Co-Exec surpasses all other baselines in aggregated throughput, it leads to a substantial degradation in training, up to 61%. Consistent trends are observed in the MP scenario shown in Figure LABEL:fig:mp-varysize. Specifically, SpecInF matches the offline throughput performance of Co-Exec while avoiding the latter’s detrimental impact on training, which can reach up to 22%. Notably, with 4 instances, the inference throughput soars to 299% more than that of Exclusive.
@@ 5.4 System Overhead
Figure 4: The system overhead of SpecInF.
@@ 6 Related Work
GPU Sharing. The straightforward method to enhance GPU utilization effectively is to share a single GPU with multiple DL tasks. Existing works can be categorized into temporal and spatial sharing approaches. For temporal sharing, Antman [24] implicitly inserts training jobs during iterations. Based on this, TGS [23] provides high transparency. However, both methods are not suitable for LLMs due to large memory footprints. As for spatial sharing, MIG [16] supports the physical isolation of GPU devices but lacks adaptability. MPS [17] is widely used in DL systems [5, 25], while it can not dynamically consume idle compute resources completely due to static allocation. Aiming at distributed training, SpecInF detects bubbles timely and speculatively fills inference workloads to improve GPU utilization.
@@ 7 Conclusion
Nowadays, Deep Learning revolutionizes various aspects of life. However, GPUs used for training these DL applications are usually underutilized, yielding massive compute and memory fragments. We observe that moderate inference workloads are well-suited to fill up these GPU fragmentations. In this paper, we present SpecInF, which collocates distributed training with online/offline inference instances, to speculatively serve inference workloads, significantly improving GPU utilization. The results show that SpecInF can exploit the idling GPU resources in various distributed training modes, delivering additional up to 14$\times$ offline inference throughputs than TGS and 67% reduction in online inference p95 latency than MPS, while guaranteeing collocated training throughput.